OpenStudy (selena2345):

Use The Distributive Property to write a sum that is equivalent to the expression . Then Find The Sum. -4(3+6+-8)

OpenStudy (anonymous):

Well do you know how to use the Distributive Property?

jimthompson5910 (jim_thompson5910):

Similar Problem -7(2+9+-11) = -7(2)-7(9)-7(-11) -7(2+9+-11) = -14-63+77 -7(2+9+-11) = 0 This isn't the answer to your problem, but it shows you how to find the answer

OpenStudy (anonymous):

-12-24+32= what

OpenStudy (selena2345):

-12-24+32=-4

OpenStudy (anonymous):

There you go!

OpenStudy (selena2345):

Thank YOU SO MUCH

OpenStudy (anonymous):

lol no problem!